complate the second schedule and pages 137 & 138
Report Form DATA Data is collected in your noteboo
Report Form DATA Data is collected in your noteboo

Report Form DATA Data is collected in your noteboo
Report Form DATA Data is collected in your noteboo
Answer #1

1.moles = mass/molar mass

mass/moles =molar mass

2. moles of any substan at STP occupies 22.4 L

moles = mass/molar mass, molar mass = moles/moles

3.

a. mole of any substance occupies 22.4 L at STP

0.033 L correspond to 0.033/22.4 L =0.001473 moles

Moles = mass/Molar mass, molar mass= mass/moles = 0.253/0.001473 g/mole =172 g/mole

b. Molar mass = 0.151/(0.029/22.4) =116 g/mole

c. molar mass = 0.023/(0.080/22.4)= 6.44g/mole

d. Molar mas = 0.131/(0.086/22.4)= 34.12 g/mole

4. g/mole is nothing but molar mass or molecular weight

Average = (172+116+6.44+34.12)/4 =82.14 gm/mole

Standard deviation =

Standard deviation = sqrt(1/4* }(172- 82.14)2+116-82.14)2 +(6.44-82.14)2 + (34.12-82.14)2 =65.68

Relative standard deviartion = 100*65.68/82.14=79.84

4. The deviations between experimental values and average values is significant. Average values is 82.14 g/mole.

5.